OpenStudy (anonymous):

Can someone explain how to multiply, divide, add, subtract fractions... I know I should know this, but fractions are always so difficult for me!

OpenStudy (anonymous):

ur supposed to flip the numerator with the denominator on the right when you dived fractions

OpenStudy (anonymous):

I definitely need more help than that haha ;(

OpenStudy (anonymous):

try khanacedemy.org

OpenStudy (anonymous):

When adding & subtracting you much first find the common denominator, then add or subtract the numerators, then reduce or simplify the fraction To multiply fractions, you will multiply the numerators then the denominators, and simplify To divide you must invert the second fraction, then follow the rules for multiplication

OpenStudy (anonymous):

\[\frac{a}{b}\div \frac{c}{d}=\frac{a}{b}\times \frac{d}{c}=\frac{a\times d}{b\times c}\] for example \[\frac{2}{3}\div \frac{5}{7}=\frac{2}{3}\times \frac{7}{5}=\frac{14}{15}\]

OpenStudy (anonymous):